OpenStudy (anonymous):

dy/dx=(ycos(x))/(1+y^2)

OpenStudy (owlcoffee):

What do you want to do with it?

OpenStudy (anonymous):

Solve for y using integration

OpenStudy (anonymous):

I know I'd get 1+y^2/y=cosx, but the integration by substitution at that point gets a bit confusing

rishavraj (rishavraj):

\[\frac{ dy }{ dx } = \frac{ y \cos(x) }{ 1 + y^2 }\] \[\int\limits\frac{ (1 + y^2)dy }{ y } = \int\limits \cos(x)dx\]

OpenStudy (owlcoffee):

Well, this is a differential equation because, we can see that a derivative appears there, so: \[\frac{ dy }{ dx }=\frac{ ycosx }{ 1+y^2 }\] so, leaving it to one side: \[\frac{ (1+y^2)dy }{ y }=cosxdx\] And now, it's a matter of integrating on both sides: \[\int\limits \frac{ (1+y^2)dy }{ y }=\int\limits cosx dx\] I will take the first steps on the left: \[\int\limits \frac{ (1+y^2)dy }{ y }\] since the grade of the numerator is higher than the denominator, we can divide both polynomials: \[\int\limits \frac{ y^2 }{ y }+\frac{ 1 }{ y }dy\] \[\int\limits\limits \frac{ y^2 }{ y }+\frac{ 1 }{ y }dy = \int\limits ydy + \int\limits \frac{ 1 }{ y }dy\] \[\int\limits \frac{ (1+y^2)dy }{ y }=\frac{ y^2 }{ 2 }+ Ln(y) +c\]

rishavraj (rishavraj):

\[\int\limits(\frac{ 1 }{ y })dy + \int\limits y dy = \sin(x) \]

OpenStudy (owlcoffee):

\[\frac{ y^2 + Ln ( y^2) }{ 2 }=sinx\]

OpenStudy (anonymous):

Okay, I see how that works Owlcoffee, thank you. When trying to solve for y what do I do when there are two ys?

OpenStudy (owlcoffee):

I'm not sure... I can't seem to be able to isolate the y...

OpenStudy (anonymous):

It's okay, I cant either

OpenStudy (owlcoffee):

Are you sure the problem is like that?

OpenStudy (anonymous):

Oh, you just have to plug in the condition given with the problem.

OpenStudy (owlcoffee):

I see, so solving wasn't needed.

OpenStudy (anonymous):

No, it was you just plug in the conditions after you get it to that point and I saw on Khan Academy that that is an infamous differential equation and you cant get any further after solving for C

OpenStudy (owlcoffee):

I'll have that in mind, thank you.

OpenStudy (anonymous):

Sure, thank you.